Choose ABCD
all fixed costs 9. If STC 300+4q +4q2 and SMC 4+8q where q is output and are sunk, the firms short-run supply curve is A) q 4+8p for p24 and zero otherwise. B) q 4+4p for p24 and zero otherwise. q o ifpcz and q-1/4p-1 if p22 D) q-o if pe4 and q-i/8p-1/2 if p24
Answer #1

Ans is D.

if p<4, q=0

If p≥4, q= 1/8p-1/2

Reason-

STC= 300+4q+4q^2

SVC= 4q+4q^2

SAVC= 4+4q

SMC=4+8q

So, the minimum of AVC (at q=0) =4

So, if p<4, q=0

If p≥4, p= 4+8q or q= 1/8p-1/2

If it helps, kindly upvote.